OpenStudy (anonymous):

The half-life of a radioactive isotope is 4 days. If 28.2 kg are present, how much will be present after 26 days?

OpenStudy (anonymous):

\[\huge 28.2\times \left(\frac{1}{2}\right)^{\frac{26}{4}}\] and a calculator

OpenStudy (anonymous):

0.31